A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red?
The probability that the first sock is blue and the second sock is red is:
(3/10) * (4/10)=3/25.
Answer:
Probability = 0.12
Step-by-step explanation:
Given : A drawer contain 4 red, 3 white , 3 blue socks
Total number of outcomes(S)= 10
Favorable event for first sock(B) ={3 blue sock}=3
Probability of picking first sock which is blue = favorable outcome/ total number of outcome P(B) = 3/10
since replacement is there
therefore, total no of outcomes remain same = 10
Favorable event for second sock(R) ={4 red sock}=4
then, probability of picking second sock which is red =favorable outcome/ total number of outcome P(R) = 4/10
Now, probability that the first sock is blue and the second sock is red is
P(B and R)= P(B) × P(R) = 3/10 ×4/10 = 12/100 = 3/25
Probability of events = 0.12